# Redefining <Creative> in Dictionary: Towards an Enhanced Semantic Understanding of Creative Generation

### v2

@@ Abstract
“Creative” remains an inherently abstract concept for both humans and diffusion models. While text-to-image (T2I) diffusion models can easily generate out-of-domain concepts like “a blue banana”, they struggle with generating combinatorial objects such as “a creative mixture that resembles a lettuce and a mantis”, due to difficulties in understanding the semantic depth of “creative”. Current methods rely heavily on synthesizing reference prompts or images to achieve a creative effect, typically requiring retraining for each unique creative output—a process that is computationally intensive and limits practical applications. To address this, we introduce CreTok, which brings meta-creativity to diffusion models by redefining “creative” as a new token, <CreTok>, thus enhancing models’ semantic understanding for combinatorial creativity. CreTok achieves such redefinition by iteratively sampling diverse text pairs from our proposed CangJie dataset to form adaptive prompts and restrictive prompts, and then optimizing the similarity between their respective text embeddings. Extensive experiments demonstrate that <CreTok> enables the universal and direct generation of combinatorial creativity across diverse concepts without additional training (4s vs. BASS’s 2400s per image), achieving state-of-the-art performance with improved text-image alignment ($\uparrow$0.03 in VQAScore) and higher human preference ratings ($\uparrow$0.009 in PickScore and $\uparrow$0.169 in ImageReward). Further evaluations with GPT-4o and user studies underscore CreTok’s strengths in advancing creative generation.
@@ 1 Introduction
Figure 1: (a) In out-of-distribution generation, diffusion models can directly generate “a blue banana” without additional training, benefiting from the clear and concrete semantics of “blue”. (b) However, they lack an intrinsic understanding of the abstract and ambiguous semantics of “creative”. (c) Leveraging the TP2O task, we redefine the token associated with “creative” as <CreTok> to bring models meta-creativity, allowing them to directly generate combinatorial creativity by enhancing their semantic understanding of “creative”.
Creativity, an inherently abstract concept even for humans [47, 27], presents a significant challenge for diffusion models. Recent efforts have sought to integrate creativity into diffusion models in more concrete ways. For instance, ConceptLab [43] approaches creativity as the generation of novel, indescribable concepts beyond conventional language. BASS [22] extends this interpretation, defining creativity as the combinatorial generation of unique objects from text pairs (e.g., (Lettuce, Mantis)), exemplified in Figure 1c. Both approaches suggest that creative ability involves generating out-of-domain images.
Diffusion models can directly generate out-of-domain images by understanding concrete prompts like “a blue banana”, but often struggle with the abstract semantics of “creative”, as illustrated in Figure 1a,b. Given the challenge that diffusion models face in directly generating creativity, existing methods typically rely on synthesizing reference prompts or images to achieve creative effects. For instance, to combine “Lettuce” and “Mantis” creatively, ConceptLab [43] merges tokens representing these concepts into a new composite token, while BASS [22] uses predefined sampling rules to search for creative outcomes from a large pool of candidate images. Similarly, personalization-based methods like MagicMix [25], DiffMorpher [56], and ATIH [51] use semantic mixing or interpolation during the diffusion process to generate novel visual representations.
@@
Through this approach, CreTok establishes <CreTok> as a universal token that brings meta-creativity to diffusion models, transforming creativity from static concept synthesis [43, 22, 51] to a more adaptable and flexible creative capability. This meta-creativity enables the model to generate novel combinatorial concepts, even when the corresponding text pairs have not been encountered during training. For instance, the combination of (Lettuce, Mantis) in Figure 1c, though unseen during training, can be creatively generated using <CreTok>. Furthermore, this meta-creativity enables direct concept combinations without requiring additional training, much like generating “a blue banana”. This significantly reduces both time and computational complexity compared to state-of-the-art (SOTA) creative generation methods, such as ConceptLab [43] (4s vs. 120s per image, 30$\times$ speedup) and BASS [22] (4s vs. 2400s per image, 600$\times$ speedup), while maintaining linguistic flexibility for diverse applications and styles.
@@
Our contributions are as follows: (1) We propose CreTok, a method designed to enhance models’ meta-ability by enabling a enhanced understanding of abstract and ambiguous adjectives (e.g., “creative” or “beautiful”) through their redefinition as new tokens. (2) Leveraging CreTok, we redefine the abstract term “creative” within our proposed CangJie dataset for the TP2O task, and introduce <CreTok>, a universal token that imparts meta-creativity to diffusion models, enabling direct application to the creative generation of diverse combinatorial concepts. (3) Experimental results demonstrate the effectiveness of CreTok in generating combinatorial creativity, outperforming SOTA text-to-image (T2I) models and creative generation methods in terms of computational complexity, human preference ratings, text-image alignment, and other key metrics.
2 Related Works
@@ 3.1 Preliminary
Latent Diffusion Models (LDMs) [44] have been widely adopted in T2I generation [13, 36, 31, 3, 54, 55]. LDMs shift the diffusion process from pixel space to a compact latent space with an encoder $\mathcal{E}$, which maps images $x$ into spatial latent codes $z=\mathcal{E}(x)$. The diffusion model is then trained to generate these latent codes through denoising, minimizing the following objective:
@@
where $\varepsilon_{\theta}$ represents the noise prediction network, trained to estimate the noise $\varepsilon$ added to the latent variable $z_{t}$ at timestep $t$, conditioned on a vector $c_{\theta}(y)$, which maps the input $y$ (e.g., text prompt) to a condition vector.
In our study, we utilize Stable Diffusion 3 (SD 3) [10] as the T2I model, which employs a transformer-based architecture to facilitate bidirectional information flow between image and text tokens. Our work focuses on the text encoder in $c_{\theta}$ and condition input $y$, while keeping the parameters of other components frozen (See Appendix 8.1 for details).
@@ 3.2 Creative Generation from a Single Text Pair
Given a text pair $(t_{1},t_{2})$ (e.g., (Lettuce, Mantis)), we generate a restrictive prompt $\mathcal{P}_{r}(t_{1},t_{2})$ by combining the pair into a phrase like “a $t_{1}$ $t_{2}$” (e.g., “a lettuce mantis.”). The trainable token <CreTok>, which redefines “creative”, is then used to form an adaptive prompt $\mathcal{P}_{a}$ representing the combinatorial results (e.g., “a photo of a <CreTok> mixture.”). To optimize the semantic alignment of $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$, we increase the similarity between the embeddings of $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$ using the following objective:
@@
As highlighted in ConceptLab [43], models may increase such similarity between $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$ by generating two distinct concepts independently (i.e., a lettuce and a mantis) rather than combining them into a unified concept. To address this, we establish a loss threshold $\theta$ to mitigate overfitting to individual concepts and reduce excessive semantic carryover (see Section 6.2 for detailed analysis).
~$\tilde{\mathcal{L}}_{\text{mix}}=1-\max[\text{cos}(E(\mathcal{P}_{r}(t_{1},t_{2})),E(\mathcal{P}_{a})),\theta]$ (3)
@@ 3.3 Refining <CreTok> in a Continuous Process
To achieve such meta-creativity, we construct a dataset specifically for the TP2O task, termed CangJie, comprising diverse text pairs (see Appendix 9 for details). Then we iteratively refine <CreTok> on CangJie through a continuous training process, gradually embedding generalized semantics of “creative” into <CreTok>. In each training iteration, a set of $n$ text pairs is randomly sampled, and the cumulative loss is calculated as:
@@
where $\mathcal{L}_{\text{mix}}$ represents the cosine similarity loss between $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$, as defined in Eq. 2. After each update, new sets of $n$ text pairs are sampled, ensuring that <CreTok> remains generalizable across a wide range of concepts.
@@ 4.1 Datasets
To comprehensively evaluate creativity, we develop CangJie, the first dataset specifically designed for the TP2O task. CangJie combines concepts from categories like animals and plants, forming text pairs through diverse combinations. The dataset includes 200 text pairs for training <CreTok>, and 27 text pairs from the original BASS [22] results for unified comparison. Detailed specifications are provided in Appendix 9.
@@ 4.2 Experimental Setup
Our implementation is based on the official Stable Diffusion 3 [10], which integrates three text encoders: CLIP L/14 [39], OpenCLIP bigG/14 [6], and T5-v1.1-XXL [40]. In our experiments, only the two CLIP models are used as text encoders $E(\cdot)$ without significant performance loss, owing to the simplicity of prompts. The training of <CreTok> runs for 10K steps on a single NVIDIA 4090 GPU, using an initial learning rate of 0.01 with a cosine scheduler, a batch size of 1, and a sample size $n=16$. The training process can be completed within approximately 30 minutes. Notably, there is NO additional computational overhead after the training of <CreTok>.
@@ 5.1.1 Comparison with State-of-the-Art T2I Models
In contrast, CreTok significantly enhances the model’s semantic understanding of “creative” through <CreTok>, enabling a more cohesive and realistic fusion while preserving the interpretability of each component. Additional images generated by CreTok are available in Appendix 11.
Figure 4: Visual comparisons of combinatorial creativity. We compare CreTok with BASS [22], and other methods achieving similar combinatorial effects, including MagicMix [25], Black-Scholes [20] and Diffmorpher [56], to highlight CreTok’s superior performance. For fair comparison, most images from these methods are sourced directly from the original papers, with a white watermark added in the bottom right corner. Additionally, generation time per image is recorded to emphasize CreTok’s meta-creativity and zero-shot capability.
@@ 5.1.2 Comparison with Creative Generation Methods
Beyond comparisons with SOTA T2I models, we evaluate CreTok against methods specifically designed for creativity and personalization to further highlight its advantages. Creative generation methods like BASS [22] achieve creative outputs through rule-based searches across large-scale candidate generations, while personalization methods, such as MagicMix [25], Black Scholes [20] and Diffmorpher [56], can generate combination effects via interpolation between noise, prompts and LoRA [17].
As shown in Figure 4, interpolation techniques used in personalization can produce similar visual effects, but they are heavily dependent on reference images, limiting their further adaptability. When significant visual disparities exist between images, the resulting fusion often lacks coherence. While BASS is capable of producing high-quality creative images without reference images, it demands substantial computational resources (2400 seconds vs. CreTok’s 4 seconds) for generation and filtering.
@@ 5.2 Creativity beyond Text Pair
5.3 Universality of <CreTok> Among Styles
Figure 6: Redefined <CreTok> can be combined with natural language to showcase combinatorial creativity in various styles. Additional styles are illustrated in Appendix 12.
5.4 Evaluation for Creativity
5.4.1 Quantitative Comparisons
@@
5.4.2 Evaluation via GPT-4o
Since existing metrics are insufficient for assessing such abstract “creativity”, we employ GPT-4o to objectively assess image creativity through quantitative analysis across four dimensions: Integration, Alignment, Originality, and Aesthetics. Detailed prompts are available in Appendix 10.
@@ 5.4.3 User Study
To comprehensively evaluate creativity, we conduct a user study involving 50 highly educated participants. Each participant ranks the creativity of images generated by CreTok in comparison with other methods. The average ranks, summarized in Table 3, reveal that CreTok significantly outperforms the current T2I diffusion models lacking specialized design for creativity and receives a higher ranking than BASS [22], achieving an average ranking of 1.9. Further details are provided in Appendix 13.
@@ 7 Conclusion
Supplementary Material
8 More Experimental Details
8.1 Stable Diffusion 3
@@
9 More Details on Proposed CangJie Dataset
We introduce CangJie, the first dataset specifically designed for the combinatorial creativity proposed in the TP2O task. Named after 仓颉, the creator of Chinese characters, CangJie symbolizes the dataset’s focus on generating novel concepts. CangJie comprises 200 text pairs, combining common animals (e.g., dogs, cats, pigs) and plants (e.g., bananas, pineapples, lettuce). Detailed composition of the dataset are presented in Table 4.
Table 4: Deatils of CangJie.

### v3

@@ Abstract
“Creative” remains an inherently abstract concept for both humans and diffusion models. While text-to-image (T2I) diffusion models can easily generate out-of-distribution concepts like “a blue banana”, they struggle with generating combinatorial objects such as “a creative mixture that resembles a lettuce and a mantis”, due to difficulties in understanding the semantic depth of “creative”. Current methods rely heavily on synthesizing reference prompts or images to achieve a creative effect, typically requiring retraining for each unique creative output—a process that is computationally intensive and limits practical applications. To address this, we introduce CreTok, which brings meta-creativity to diffusion models by redefining “creative” as a new token, <CreTok>, thus enhancing models’ semantic understanding for combinatorial creativity. CreTok achieves such redefinition by iteratively sampling diverse text pairs from our proposed CangJie dataset to form adaptive prompts and restrictive prompts, and then optimizing the similarity between their respective text embeddings. Extensive experiments demonstrate that <CreTok> enables the universal and direct generation of combinatorial creativity across diverse concepts without additional training, achieving state-of-the-art performance with improved text-image alignment and higher human preference ratings. Code will be made available at https://github.com/fu-feng/CreTok.
@@ 1 Introduction
Figure 1: (a) In out-of-distribution generation, diffusion models can directly generate “a blue banana” without additional training, benefiting from the clear and concrete semantics of “blue”. (b) However, they lack an intrinsic understanding of the abstract and ambiguous semantics of “creative”. (c) Leveraging the TP2O (i.e., Creative Text Pair to Object) task, we redefine the token associated with “creative” as <CreTok> to bring models meta-creativity, allowing them to directly generate combinatorial creativity by enhancing their semantic understanding of “creative”.
Creativity, an inherently abstract concept even for humans [48, 28], presents a significant challenge for diffusion models. Recent efforts have sought to integrate creativity into diffusion models in more concrete ways. For instance, ConceptLab [44] approaches creativity as the generation of novel, indescribable concepts beyond conventional language. BASS [23] extends this interpretation, defining creativity as the combinatorial generation of unique objects from text pairs (e.g., (Lettuce, Mantis)), exemplified in Figure 1c. Both approaches suggest that creative ability involves generating out-of-distribution images.
Diffusion models can directly generate out-of-distribution images by understanding concrete prompts like “a blue banana”, but often struggle with the abstract semantics of “creative”, as illustrated in Figure 1a,b. Given the challenge that diffusion models face in directly generating creativity, existing methods typically rely on synthesizing reference prompts or images to achieve creative effects. For instance, to combine “Lettuce” and “Mantis” creatively, ConceptLab [44] merges tokens representing these concepts into a new composite token, while BASS [23] uses predefined sampling rules to search for creative outcomes from a large pool of candidate images. Similarly, personalization-based methods like MagicMix [26], DiffMorpher [57], and ATIH [52] use semantic mixing or interpolation during the diffusion process to generate novel visual representations.
@@
Through this approach, CreTok establishes <CreTok> as a universal token that brings meta-creativity to diffusion models, transforming creativity from static concept synthesis [44, 23, 52] to a more adaptable and flexible creative capability. This meta-creativity enables the model to generate novel combinatorial concepts, even when the corresponding text pairs have not been encountered during training. For instance, the combination of (Lettuce, Mantis) in Figure 1c, though unseen during training, can be creatively generated using <CreTok>. Furthermore, this meta-creativity enables direct concept combinations without requiring additional training, much like generating “a blue banana”. This significantly reduces both time and computational complexity compared to state-of-the-art (SOTA) creative generation methods, such as ConceptLab [44] (4s vs. 120s per image, 30$\times$ speedup) and BASS [23] (4s vs. 40s per image, 10$\times$ speedup), while maintaining linguistic flexibility for diverse applications and styles.
@@
Our contributions are as follows: (1) We propose CreTok, a method designed to enhance models’ meta-ability by enabling an enhanced understanding of abstract and ambiguous adjectives (e.g., “creative” or “beautiful”) through their redefinition as new tokens. (2) Leveraging CreTok, we redefine the abstract term “creative” within our proposed CangJie dataset for the TP2O task, and introduce <CreTok>, a universal token that imparts meta-creativity to diffusion models, enabling direct application to the creative generation of diverse combinatorial concepts. (3) Experimental results demonstrate the effectiveness of CreTok in generating combinatorial creativity, outperforming SOTA text-to-image (T2I) models and creative generation methods in terms of computational complexity, human preference ratings, text-image alignment, and other key metrics.
2 Related Work
@@ 3.1 Preliminary
Latent Diffusion Models (LDMs) [45] have been widely adopted in T2I generation [13, 37, 32, 3, 55, 56, 17, 18]. LDMs shift the diffusion process from pixel space to a compact latent space with an encoder $\mathcal{E}$, which maps images $x$ into spatial latent codes $z=\mathcal{E}(x)$. The diffusion model is then trained to generate these latent codes through denoising, minimizing the following objective:
@@
where $\varepsilon_{\theta}$ represents the noise prediction network, trained to estimate the noise $\varepsilon$ added to the latent variable $z_{t}$ at timestep $t$, conditioned on $c_{\theta}(y)$, a vector derived from the input $y$ (e.g., text prompt) through a mapping function $c_{\theta}$.
In our study, we utilize Stable Diffusion 3 (SD 3) [10] as the base model, which employs a transformer-based architecture to facilitate bidirectional information flow between image and text tokens. Our work focuses on the text encoder in $c_{\theta}$ and condition input $y$, while keeping the parameters of other components frozen (See Appendix A for details).
@@ 3.2 Creative Generation from a Single Text Pair
As shown in Figure 2, given a text pair $(t_{1},t_{2})$ (e.g., (Lettuce, Mantis)), we generate a restrictive prompt $\mathcal{P}_{r}(t_{1},t_{2})$ by combining the pair into a phrase like “a $t_{1}$ $t_{2}$” (e.g., “a lettuce mantis.”). The trainable token <CreTok>, which redefines “creative”, is then used to form an adaptive prompt $\mathcal{P}_{a}$ representing the combinatorial results (e.g., “a photo of a <CreTok> mixture.”). To optimize the semantic alignment of $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$, we increase the similarity between the embeddings of $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$ using the following objective:
@@
As noted in ConceptLab [44], overfitting can artificially inflate similarity by disproportionately reinforcing one concept while neglecting others. To address this issue, we introduce a loss threshold $\theta$ to regulate concept integration. Moreover, to ensure the coherent fusion of two concepts, rather than their independent generation (e.g., a lettuce and a mantis), $\theta$ must remain moderate to avoid low similarity between $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$ (see Section 6.2 for details).
~$\tilde{\mathcal{L}}_{\text{mix}}=1-\min[\text{cos}(E(\mathcal{P}_{r}(t_{1},t_{2})),E(\mathcal{P}_{a})),\theta]$ (3)
@@ 3.3 Refining <CreTok> in a Continuous Process
To achieve such meta-creativity, we construct a dataset specifically for the TP2O task, termed CangJie, comprising diverse text pairs (see Appendix D for details). Then we iteratively refine <CreTok> on CangJie through a continuous training process, gradually embedding generalized semantics of “creative” into <CreTok>. In each training iteration, a set of $n$ text pairs is randomly sampled, and the cumulative loss is calculated as:
@@
where $\mathcal{L}_{\text{mix}}$ represents the cosine similarity loss between $\mathcal{P}_{r}$ and $\mathcal{P}_{a}$, as defined in Eq. (2). After each update, new sets of $n$ text pairs are sampled, ensuring that <CreTok> remains generalizable across a wide range of concepts.
@@ 4.1 Datasets
To comprehensively evaluate creativity, we develop CangJie, the first dataset specifically designed for the TP2O task. CangJie combines concepts from categories like animals and plants, forming text pairs through diverse combinations. The dataset includes 200 text pairs for training <CreTok>, and 27 text pairs from the original BASS [23] results for unified comparison. Detailed specifications are provided in Appendix D.
@@ 4.2 Experimental Setup
Our implementation is based on the official Stable Diffusion 3 [10], which integrates three text encoders: CLIP L/14 [40], OpenCLIP bigG/14 [6], and T5-v1.1-XXL [41]. In our experiments, only the two CLIP models are used as text encoders $E(\cdot)$ without significant performance loss, owing to the simplicity of prompts. The training of <CreTok> runs for 10K steps on a single NVIDIA 4090 GPU, using an initial learning rate of 0.01 with a cosine scheduler, a batch size of 1, and gradient accumulation over $n=16$ steps. The training process can be completed within approximately 30 minutes. Notably, there is NO additional computational overhead after the training of <CreTok>.
@@ 4.3 Evaluation Metrics
Figure 4: Visual comparisons of combinatorial creativity. We compare CreTok with BASS [23], and other methods achieving similar combinatorial effects, including MagicMix [26] and Black-Scholes [21], to highlight CreTok’s superior performance. For fair comparison, most images from these methods are sourced directly from the original papers, with a white watermark added in the bottom right corner. Additionally, generation time per image is recorded to emphasize CreTok’s meta-creativity and zero-shot capability.
@@ 5.1.1 Comparison with State-of-the-Art T2I Models
In contrast, CreTok significantly enhances the model’s semantic understanding of “creative” through <CreTok>, enabling a more cohesive and realistic fusion while preserving the interpretability of each component. Additional images generated by CreTok are available in Appendix B.2.
@@ 5.1.2 Comparison with Creative Generation Methods
Beyond comparisons with SOTA T2I models, we evaluate CreTok against methods specifically designed for creativity and personalization to further highlight its advantages. Creative generation methods like BASS [23] achieve creative outputs through rule-based searches across large-scale candidate generations, while personalization methods, such as MagicMix [26] and Black Scholes [21], can generate combination effects via interpolation between noise, prompts.
As shown in Figure 4, interpolation techniques used in personalization can produce similar visual effects, but they are heavily dependent on reference images, limiting their further adaptability. When significant visual disparities exist between images, the resulting fusion often lacks coherence (see Appendix B.1 for additional comparisons). While BASS is capable of producing high-quality creative images without reference images, it demands substantial computational resources (40 seconds vs. CreTok’s 4 seconds) for generation and filtering.
@@ 5.2 Creativity beyond Text Pair
Figure 6: Redefined <CreTok> can be combined with natural language to showcase combinatorial creativity in various styles. Additional styles are illustrated in Appendix B.3.
@@
5.3 Universality of <CreTok> Among Styles
5.4 Evaluation for Creativity
5.4.1 Quantitative Comparisons
5.4.2 Evaluation via GPT-4o
Since existing metrics are insufficient for assessing such abstract “creativity”, we employ GPT-4o to objectively assess image creativity through quantitative analysis across four dimensions: Integration, Alignment, Originality, and Aesthetics. Detailed prompts are available in Appendix C.1.
@@ 5.4.3 User Study
To comprehensively evaluate creativity, we conduct a user study involving 50 highly educated participants. Each participant ranks the creativity of images generated by CreTok in comparison with other methods. The average ranks, summarized in Table 3, reveal that CreTok significantly outperforms the current T2I diffusion models lacking specialized design for creativity and receives a higher ranking than BASS [23], achieving an average ranking of 1.9. Further details are provided in Appendix C.2.
@@ 7 Conclusion
Acknowledgement
We sincerely thank Wenqian Li, Jianlu Shen, and Ruixiao Shi for their insightful discussions on this work. We also appreciate Freepik for contributing to the figure design. This research is supported by the Science and Technology Major Project of Jiangsu Province under Grant BG20240305, Key Program of Jiangsu Science Foundation under Grant BK20243012, National Natural Science Foundation of China under Grants U24A20324, 62125602, and 62306073, Natural Science Foundation of Jiangsu Province under Grant BK20230832, and the Xplorer Prize.
Appendix A More Details on Stable Diffusion 3
@@
Appendix B Additional Results
Figure B.1: More visual comparisons of combinatorial creativity. We further compare CreTok with MagicMix [26] and DiffMorpher [57], to further highlight CreTok’s superior performance.
B.1 Additional Comparison Results
We further compare CreTok with other personalization methods designed to produce combination effects through interpolation, such as MagicMix [26] and Diffmorpher [57].
As illustrated in Figure B.1, interpolation-based techniques heavily rely on reference images, which constrains their adaptability. When substantial visual differences exist between the input images, the resulting fusion often appears incoherent, a limitation particularly evident in Diffmorpher.
In contrast, CreTok enables the diffusion model to directly generate combinatorial creativity without relying on image synthesis, producing outputs with enhanced visual coherence and detail.
B.2 Additional Creative Visual Results
We present additional creative images generated by CreTok in Figure B.2, further demonstrating the universality of <CreTok> in imparting meta-creativity to diffusion models. For each text pair ($t_{1}$, $t_{2}$), the corresponding combinatorial creativity is produced using the prompts: “A photo of a <CreTok> mixture that resembles $t_{1}$ and $t_{2}$”.
Figure B.2: Additional results of combinatorial creativity generated by CreTok.
B.3 Additional Styles of Creativity
In addition to the creative image styles presented in Figure 6, Figure B.3 showcases additional results, further illustrating the universality of <CreTok>. These results further highlight its seamless integration with natural language, enabling diverse and flexible concept combinations across various styles.
Figure B.3: More styles of combinatorial creativity generated by CreTok.
Appendix C More Details on Evaluation
C.1 Prompts Used for GPT-4o Evaluation
C.2 More Details on User Study
This section provides a detailed overview of our User Study. The interface used for the study is illustrated in Figure C.4. We utilize 27 text pairs sourced from the original BASS paper to ensure a fair comparison. Creative images generated by CreTok and other methods are displayed in Figure C.1, C.2, and C.3. This user study involves 50 participants who evaluate and rank the creative outputs for each text pair (1-5). Their responses are summarized in Table C.1.
Figure C.1: Images generated by CreTok and other methods used in the User Study.
Figure C.2: Images generated by CreTok and other methods used in the User Study (continued).
Figure C.3: Images generated by CreTok and other methods used in the User Study (continued).
Figure C.4: Interface of the User Study.
Table C.1: Average ranking of each text pair evaluated in the User Study.
#1 #2 #3 #4 #5 #6 #7 #8 #9 #10 #11 #12 #13 #14
Stable Diffusion 3 1.86 2.50 2.74 4.04 3.74 2.26 2.22 3.04 4.58 2.62 2.98 3.66 2.54 2.84
Stable Diffusion 3.5 3.52 4.20 3.76 1.98 3.02 3.02 2.48 2.96 2.88 3.50 3.00 2.88 2.90 3.60
Kandinsky 3 3.90 2.64 3.40 2.70 2.68 3.00 3.78 2.26 1.70 2.86 3.30 2.14 3.82 2.54
BASS 3.90 3.82 2.98 3.52 2.08 3.76 2.86 4.38 3.44 3.70 2.98 3.64 2.92 3.44
CreTok 1.82 1.84 2.12 2.76 3.48 2.96 3.66 2.36 2.40 2.32 2.74 2.68 2.82 2.58
#15 #16 #17 #18 #19 #20 #21 #22 #23 #24 #25 #26 #27
Stable Diffusion 3 4.28 3.16 3.00 3.78 2.92 4.20 3.40 2.54 2.56 4.64 4.18 3.78 4.36
Stable Diffusion 3.5 3.90 1.88 3.00 3.32 2.56 2.12 3.22 2.30 2.76 3.04 3.62 1.96 3.04
Kandinsky 3 3.08 3.70 3.58 3.50 3.98 3.32 4.14 3.66 3.40 2.52 2.96 4.28 2.68
BASS 1.56 3.14 2.64 2.14 3.44 3.38 2.20 2.74 3.74 2.62 2.94 3.28 2.20
CreTok 2.18 3.12 2.78 2.26 2.10 1.98 2.04 3.76 2.54 2.18 1.30 1.70 2.72
Appendix D More Details on Proposed CangJie Dataset
We introduce CangJie, the first dataset specifically designed for the combinatorial creativity proposed in the TP2O task. Named after 仓颉, the creator of Chinese characters, CangJie symbolizes the dataset’s focus on generating novel concepts. CangJie comprises 200 text pairs, combining common animals (e.g., dogs, cats, pigs) and plants (e.g., bananas, pineapples, lettuce). Detailed composition of the dataset are presented in Table D.1.
Table D.1: Deatils of CangJie.